TMS Meeting Minutes, March 11, 2021

Attendees: Andrew, Breann, Heejoo, Olga, Trenton (Zoom Meeting)


Updates from the past week:
  • John tested the updated TMS code (all good)
  • Ignored guiding offset, able to correctly compensate global offset
  • Yang updated the std. dev. filter with a 5 micron threshold
  • Measured the time of the TMS measurement cycle
    • 1 measurement set: 8-9 seconds
    • TMS measurement + PSF correction + overhead: 20 seconds
    • Ref. Vec. Cycle (3 TMS cycles + PSF + overhead): 39 seconds
  • Seems we have a system that is functionally ready
  • Heejoo installed the new beta version of the Etalon software
  • All systems appear to be working well right now
  • Heejoo keeping an eye on the laser power monitor to ensure there is no leakage
LBT Visit Next Week:
  • Heejoo visiting next Tuesday to Thursday
  • Planning to finish installing system on the SX side
  • Stretch goal to achieve the laser tracker survey, if able (Mike Gardner won't be on the mountain but planning to help remotely)
  • Andrew suggests helping remotely to drive the laser tracker (planning to prepare some automation in advance)
    • Will have to move the laser tracker once during measurement
    • Andrew suggested a method using measurements of all collimators from both reference positions
  • Most likely able to start the laser tracker survey on Thursday morning
  • Heejoo hoping to finish all of the SX side installation next week
Technical Nights:
  • Towards the end of Saturday night is available
  • Penciled in EMS testing with LUCI, then able to switch to LBC around 3:00 AM
  • Next step is to run Yang's code in control mode for a long time (probe for bugs)
  • Observe for a long time in a field with a large change (e.g. transit going through the meridian)
  • Obtain to defocus images with VFAST Extra analysis (more quantitative information with aberrations than with focused images)
    • Able to resolve different aberration contributions with donuts
    • Also able to obtain aberrations on both red and blue sides
    • Obtaining direct aberration residuals post TMS-correction
    • Could be quite useful to have all four chips with aberration maps
    • Also useful for obtaining a more complete picture of the total telescope misalignments
  • Primary focus is testing the TMS performance (secondarily able to pull out more information after the fact)
  • Looking for two hours of time (more would be nice), one hour at a minimum

  • Would also be good to have some passive observing nights
  • Potential run the system passively (LBC in use Wed., Thu., Fri., Sat.)
  • Heejoo might be able to help cover the operation of Yang's code along with Olga
Collimators:
  • Thorlabs is still working on evaluating the performance and cost (sending update by tomorrow)
    • Changing spacings and replacing lenses with other off the shelf lenses
    • Also capable of making a new barrel
  • Edmund Optics gave up; unable to achieve the required adjustment by fabricating a new barrel
  • Andrew requests to know what the cost comes out to (Kiwi Star could do better)
  • Heejoo currently has one of the solid retros
    • Andrew requests a quick test of the retro at two distances

-- TrentonBrendel - 11 Mar 2021
Topic revision: r1 - 11 Mar 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback